OpenStudy (anonymous):

Assume that a procedure yields a binomial distribution with a trial repeated n times. Use the binomial probability formula to find the probability of x successes given the probability p of success on a single trial. Round to three decimal places n = 4, x = 3, p =1/3

OpenStudy (aravindg):

Doyou know the formula for binomial distribution?

OpenStudy (aravindg):

\[\large P(X=r)=\left(\begin{matrix}n \\ r\end{matrix}\right)p^r q^{n-r}\]

OpenStudy (aravindg):

p+q=1. Plug in the values. Get your answer.